An alkaloid was isolated from a common household beverage. The unknown alkaloid proved to have a molecular mass of 194. Using the Rule of Thirteen, determine a molecular formula and an index of hydrogen deficiency for the unknown. Alkaloids are naturally occurring organic substances that contain nitrogen. (Hint: There are four nitrogen atoms and two oxygen atoms in the molecular formula. The unknown is caffeine. Look up the structure of this substance in The Merck Index and confirm its molecular formula.)


1
Expert's answer
2021-02-15T03:22:14-0500

The rule of 13 states that the formula of a compound is a multiple n of 13 (the molar mass of CH) plus a remainder r. From here:

194 / 13 = 14r12

The predicted formula is C14H14+12 or C14H26.

However, as the molecule contains four nitrogen atoms and two oxygen atoms, we need to substract the corresponding values:

1. 4 × [N - 15] = 4 × [N - (CH + H)] = 4 × [N - CH2].

This gives a formula C10H18N4.

2. 2 × [O - 16] = 2 × [O - CH4].

This gives a final formula C8H10N4O2.

Index of hydrogen deficiency can be calculated using the following formula:

Index = (2C + 2 + N - H) / 2 = (2 × 8 + 2 + 4 - 10) / 2 = 6.


Answer:C8H10N4O2; index 6.